1. The basic WACC equation

The calculation of WACC involves calculating the weighted average of the required rates of return on debt, preferred stock, and common equity, where the weights equal the percentage of each type of financing in the firm’s overall capital structure.

____ is the symbol that represents the cost of raising capital through retained earnings in the weighted average cost of capital (WACC) equation.

Bryant Co. has $2.7 million of debt, $1 million of preferred stock, and $2.1 million of common equity. What would be its weight on common equity?

a. 0.14

b. 0.17

c. 0.36

d. 0.47
Answer #1

a.

Rs is the symbol that represents the cost of raising capital through retained earnings in the weighted average cost of capital (WACC) equation.

b.Total value=(2.7+1+2.1)=5.8 million

Hence weight of equity=(2.1/5.8)

=0.36(Approx).